OpenStudy (anonymous):

4(y-4)=8 A)-2 B)2 C)4 D)6

rishavraj (rishavraj):

a(x - y) = ax - ay

OpenStudy (anonymous):

A

OpenStudy (anonymous):

thank you

OpenStudy (texaschic101):

I wouldn't just take someones answer until you are sure they did it correctly....or at least explained it to you.

rishavraj (rishavraj):

@Shelby585849 its not A.... and @DALLINATOR720 dont give direct answers

OpenStudy (anonymous):

y+4 8+4 y 12 y=12?

rishavraj (rishavraj):

first of all expand it ...as i mentioned before a(x - y) = ax - ay

OpenStudy (texaschic101):

your forgetting about the 4 on the outside of the parenthesis Shelby

OpenStudy (anonymous):

4/(12-4) 4/8 y=2?

rishavraj (rishavraj):

4(y - 4) = 8 4y - 16 = 8 proceed .....

OpenStudy (texaschic101):

Example : 3(x + 2).....this is the same as saying 3 * x + 3 * 2 = 3x + 6 your basically multiplying the number on the outside of the parenthesis, with everything on the inside

OpenStudy (texaschic101):

now that you have gotten rid of the parenthesis... 4y - 16 = 8......try adding 16 to both sides

OpenStudy (anonymous):

8+16=24 4 + 16=20?

OpenStudy (texaschic101):